### v1

# Symbolic recursion method for strongly correlated fermions in two and three dimensions

Abstract
We present a symbolic implementation of recursion method for the dynamics of strongly correlated fermions on one-, two- and three-dimensional lattices. Focusing on two paradigmatic models, interacting spinless fermions and the Hubbard model, we first directly confirm that the universal operator growth hypothesis holds for interacting fermionic systems, manifested by the linear growth of Lanczos coefficients. Equipped with symbolically computed Lanczos coefficients and knowledge of their asymptotics, we are able to compute infinite-temperature autocorrelation functions up to times long enough for thermalization to occur. In turn, the knowledge of autocorrelation functions unlocks transport properties. We compute with high precision the charge diffusion constant over a broad range of interaction strengths, $V$. Surprisingly, we observe that these results are well described by a simple functional dependence $\sim 1/V^{2}$ universally valid both for small and large $V$. All results are obtained directly in the thermodynamic limit. Our results highlight the promise of symbolic computational paradigm where the most costly step is performed once and outputs symbolic results that can further be used multiple times to easily compute physical quantities for specific values of model parameters.
Dynamics and transport in strongly correlated electron systems, such as the spinless-fermion $t$-$V$ model and the Hubbard model, have been at the forefront of condensed-matter physics for decades [laughlin1983anomalous, dagotto1994correlated, imada1998metal, tokura2003correlated, troyer2005computational, brown2019bad, nichols2019spin, guardado2020subdiffusion, kovavcevic2025toward]. Accurate simulation of these models remains one of the central challenges of modern theoretical physics. Approaches based on matrix diagonalization are fundamentally limited by the exponential growth of the many-body Hilbert space [yamada200516, innerberger2020electron]. In one dimension, remarkable progress has been achieved using integrability-based techniques [prosen2011exact, bertini2021finite] and tensor-network methods [Wall_2012], to the point that a broad class of dynamical and transport problems, with certain reservations, can now be considered simulable in 1D. In higher dimensions the substantial progress has been achieved using a variety of analytical and numerical approaches [barthel2009contraction, kraus2010fermionic, corboz2010simulation, pivzorn2010fermionic, bergeron2011optical, Shakirov_2015_Modeling, perepelitsky2016transport, vucicevic2018conductivity, kiely2021transport, ulaga2022thermal, ulaga2023thermoelectric, vuvcivcevic2023charge, kovavcevic2025toward]. In parallel, significant efforts have been directed toward quantum simulation platforms, which aim to synthetically realize correlated many-body systems and probe their properties experimentally [schreiber2015observation, arute2010observation, vilchez2025extracting, chowdhury2025quantum, alam2025fermionic]. Despite impressive recent progress, such platforms remain at an early stage of development and would greatly benefit from reliable computational benchmarks capable of guiding both experiments and theory.
Figure 1: Conceptual scheme of the symbolic–computational workflow: the computationally expensive symbolic evaluation of the $n$-fold commutators $H(V),A]^{(n)}$ is carried out once, while the resulting symbolic moments $\mu_{2n}(V)$ are subsequently reused multiple times to compute correlation functions $C(t)$ at specific values of the parameter $V$.
In this work, we develop a symbolic computational approach to the dynamics of strongly correlated fermionic systems based on the recursion method [Mori_1965_Continued-fraction, Dupuis_1967_Moment, haydock1980recursive, viswanath1994recursion, mattis1981reduce]. While the method was originally introduced decades ago, recent advances in the theoretical understanding of operator growth [Elsayed_2014_Signatures, Bouch_2015_Complex, parker2019universal, Avdoshkin_2020_Euclidean, De_2024_Stochastic, Teretenkov_2025_Pseudomode, nandy2025quantum, Gamayun_2025_Exactly, loizeau2025opening, Pinna_2025_Approximation] and the increase of computational capabilities have renewed interest in the recursion method and other operator-based approaches to quantum dynamics. The recursion method has been successfully applied in a variety of settings, including interacting spin-$1/2$ systems [viswanath1994ordering, yates2020lifetime, yates2020dynamics, Uskov_2024_Quantum, shirokov2025quench] and higher-spin models [ermakov2025operator]. Other Heisenberg evolution based methods have also been developed [Rakovszky_2022_Dissipation-assisted, ermakov2024unified, schuster2025polynomial, Begusic_2025_Real-time, angrisani2025simulating, rudolph2025pauli, loizeau2025quantum], predominantly formulated in qubit operator spaces where the Pauli-string basis provides a natural framework.
Fermionic models pose additional challenges, due to anticommutation relations of fermionic operators on different sites. Recently, a symbolic Majorana-propagation algorithm for fermionic circuits has been introduced in Ref. [miller2025simulation] generalizing similar algorithms for qubit systems [schuster2025polynomial, ermakov2024unified, rudolph2025pauli, Begusic_2025_Real-time].
We report the successful application of the advanced recursion method to interacting fermionic systems on one-, two-, and three-dimensional lattices. In one dimension, the recursion method is known to be non-competitive with state-of-the-art approaches based on exact diagonalization or matrix product states [Yi-Thomas_2024_Comparing, Uskov_2024_Quantum, shirokov2025quench]. We therefore employ it in one dimension primarily for the purpose of cross-checking our results against exact diagonalization. In contrast, real-time dynamics in two and three spatial dimensions is widely believed to be challenging for established numerical methods. We establish the advanced recursion method as a powerful technique applicable in higher dimensions, capable of capturing the dynamics over the entire time domain – from short times to thermalization – in the strongly correlated regime and without finite-size effects.
The rest of the manuscript is organized as follows. In the next section, we introduce the models under study. We then discuss the autocorrelation function and its moments. After that, we briefly review the recursion method and the universal operator growth hypothesis (UOGH) [parker2019universal], and demonstrate that our results confirm the validity of the UOGH for interacting fermions in one and two dimensions. We subsequently present results for infinite-temperature autocorrelation functions obtained using the recursion method supplemented by the UOGH. We then apply the recursion method to describe transport in the models under study. Specifically, we compute the diffusion constant and the conductivity for a broad range of interaction strengths. Finally, we describe important details of the algorithm and software implementation in a separate section. The manuscript concludes with a summary and outlook.
Models and observables
We consider two fermionic models on hypercubic lattices. The first is the spinless fermion $t$-$V$ model
$H_{\mathrm{tV}}$ $=-\,t_{\text{hop}}\sum_{\langle i,j\rangle}\left(c_{i}^{\dagger}c_{j}+\text{H.c.}\right)+V\sum_{\langle i,j\rangle}\Bigl(n_{i}-\tfrac{1}{2}\Bigr)\Bigl(n_{j}-\tfrac{1}{2}\Bigr).$ (1)
The second is the Hubbard model,
$H_{\mathrm{Hub}}$ $=-t_{\text{hop}}\sum_{\langle i,j\rangle,\sigma}\left(c_{i\sigma}^{\dagger}c_{j\sigma}+\text{H.c.}\right)$
$~~~+V\sum_{i}\left(n_{i\uparrow}-\tfrac{1}{2}\right)\left(n_{i\downarrow}-\tfrac{1}{2}\right).$ (2)
Here $c^{\dagger}_{i\sigma}$ and $c_{i\sigma}$ are fermionic creation and annihilation operators for spin $\sigma=\uparrow,\downarrow$ on site $i$, satisfying $\{c_{i\sigma},c_{j\tau}^{\dagger}\}=\delta_{ij}\,\delta_{\sigma\tau}$. Throughout the text, whenever the spin index is omitted, it refers to the spinless case. The number operator is $n_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}$. The parameters $t_{\text{hop}}$ and $V$ denote the hopping amplitude and interaction strength, we set $t_{\text{hop}}=1$ unless stated otherwise. $\langle i,j\rangle$ labels nearest–neighbour sites. We adopt the convention $\hbar=1$ and $k_{\rm B}=1$ throughout the paper.
We choose the total current along a fixed lattice direction, $\hat{x}$, as an observable of interest. For the Hubbard model the current reads
$J=-\,i\sum_{i,\sigma}\left(c_{i\sigma}^{\dagger}c_{i+\hat{x},\sigma}-c_{i+\hat{x},\sigma}^{\dagger}c_{i\sigma}\right),$ (3)
where $i+\hat{x}$ denotes the nearest neighbour of site $i$ displaced by one lattice spacing along the chosen direction (in one dimension this reduces to $i+1$). For the $t$-$V$ model the current is given by the same formula with subscripts $\sigma$ dropped.
Figure 2: Dynamics of infinite-temperature autocorrelation function (4) computed using the recursion method. The results are valid in thermodynamic limit. Top row: spinless-fermion $t$-$V$ model (1). Bottom row: Hubbard model (2). Curves of different colors correspond to different values of the interaction parameter $V=2,4$, while being obtained from the same set of symbolically computed moments. Gray dashed lines indicate rigorous upper and lower bounds (7) obtained from the Taylor expansion. Green dots are the exact diagonalization results for a periodic chains of $L=12$ and $L=7$ sites in panels (a) and (d) correspondingly. Insets show the Lanczos coefficients and their linear extrapolation.
Autocorrelation function and moments.
We introduce an inner product in the space of operators according to $(A|B)\equiv\mathrm{tr}(A^{\dagger}B)/\mathrm{dim}(\mathcal{H})$, where $A$ and $B$ are arbitrary operators and $\mathrm{dim}(\mathcal{H})$ is the (finite) Hilbert space dimension. The corresponding operator norm reads $\lVert A\rVert=\sqrt{(A|A)}$.
We wish to compute the infinite-temperature autocorrelation function of the current defined as
$C(t)\equiv\frac{(J(t)|J)}{\lVert J\rVert^{2}}.$ (4)
Here $J(t)=e^{it\mathcal{L}}J$ is the current operator (3) in the Heisenberg representation, with $\mathcal{L}\equiv[H,\cdot]$ being the Liouvillian superoperator acting on the space of operators.
Note that since the models we consider conserve the total number of fermions (in the Hubbard model – separately for each polarization), the equilibrium state of the system is characterized by average density of fermions, in addition to temperature. The above definition of the autocorrelation function corresponds to half-filling, $\langle n_{j}\rangle=1/2$ ($\langle n_{j\sigma}\rangle=1/2$ for the Hubbard model).
Moments of the autocorrelation function are defined as
$\mu_{2n}\equiv\frac{(\mathcal{L}^{2n}J|J)}{\lVert J\rVert^{2}}=\frac{(\mathcal{L}^{n}J|\mathcal{L}^{n}J)}{\lVert J\rVert^{2}}=(-1)^{n}\frac{d}{dt}C(t)\big|_{t=0}.$ (5)
Note that this definition implies $\mu_{0}=1$.
Importantly, moments are polynomials of parameters linearly entering the Hamiltonian. Specifically, for the Hamiltonians (1) and (2), the moments (5) are polynomials in $V$, $\mu_{2n}(V)=\sum_{k=1}^{n}c^{(n)}_{k}V^{2k}$. In addition, the fermionic commutation relations ensure that coefficients $c^{(n)}_{k}$ for the current operator (3) are rational numbers as long as the observable . In case of $t_{\text{hop}}\neq 1$, the moments scale as $\mu_{2n}(t_{\text{hop}},V)=(t_{\text{hop}})^{2n}\mu_{2n}(V/t_{\text{hop}})$.
The core routine of our code outputs the first $n_{\text{max}}$ moments as polynomials in the system parameters. The computation is performed directly in the thermodynamic limit. We have obtained the sets of coefficients $c^{(n)}_{k}$ [ErmakovSymbolicMoments] for the Hamiltonians (1) and (2) in one, two and three dimensions. As an illustration, the first few moments of the one-dimensional ttVV model (1) read
$2\mu_{2}(V)=V^{2},$
$2\mu_{4}(V)=5V^{2}+V^{4},$
$2\mu_{6}(V)=70V^{2}+21V^{4}+V^{6}.$ (6)
Once the first $n_{\text{max}}$ moments are computed, one can extract a rich of information about the dynamics of $C(t)$. In particular, one can obtain upper and lower bounds on the autocorrelation function [Platz_1973_Rigorous, Roldan_1986_Dynamic, Brandt_1986_High, Bohm_1992_Dynamic]. The simplest bound reads
$P_{4l\pm 2}(t)\leq C(t)\leq P_{4l}(t),$ (7)
where $l=[n_{\rm max}/2]$ and $P_{m}(t)$ are Taylor polynomials of order $m$ constructed from the moments (with odd moments vanishing). The bounds (7) are extremely tight initially but rapidly diverge at later times. The scope of these bounds is ultimately limited by the convergence radius $t^{\ast}$ of the Taylor expansion, which is finite for 2D and 3D systems [parker2019universal].
We note that, although the bounds (7) are not the main result of this work, they provide a valuable rigorous benchmark for the early-time behavior of $C(t)$, in some cases extending to relatively long times. We make the moments required to construct these bounds available in symbolic form [ErmakovSymbolicMoments], so that they can be used to benchmark other numerical methods and quantum simulations of the $t$-$V$ and Hubbard models for arbitrary model parameters.
Lanczos coefficients and UOGH.
[…]
In fact, Lanczos coefficients $b_{n}$ can be obtained directly from moments $\mu_{2n}$ [Dupuis_1967_Moment]. We employ this relation reviewed in [supplement] to symbolically compute Lanczos coefficients as functions of the interaction strength for the $t$-$V$ and Hubbard models in one, two and three dimensions, with the results illustrated in Fig 2.
An important question is the asymptotic behavior of Lanczos coefficients at large $n$. It can be proven for locally interacting systems that $b_{n}$ can not grow with $n$ faster than linear (with an additional logarithmic correction for one-dimensional systems) [parker2019universal]. The UOGH [parker2019universal] asserts that for chaotic quantum many-body systems this bound is saturated, i.e
$b_{n}=\alpha n+\gamma+o(1)\qquad{\rm as}\qquad n\rightarrow\infty$ (8)
in two and three spatial dimensions. For integrable systems the scaling is not universal: while typically it is either $\sqrt{n}$ or $O(1)$ [parker2019universal], linear scaling has also been observed [Dymarsky_2021_Krylov, Bhattacharjee_2022_Krylov, Camargo_2023_Krylov, Avdoshkin_2024_Krylov, Vasli_2024_Krylov, he2025krylov].
The predictions of UOGH has been previously confirmed in various spin-1/2 systems on one- [parker2019universal, Noh_2021, Heveling_2022_Numerically, De_2024_Stochastic], two- [Cao_2021_Statistical, Heveling_2022_Numerically, Uskov_2024_Quantum, De_2024_Stochastic] and three-dimensional [Cao_2021_Statistical, Heveling_2022_Numerically, shirokov2025quench] lattices, for higher-spin systems on one- and two-dimensional lattices [ermakov2025operator], as well as for a three-site system of interacting bosons [Bhattacharyya_2023_Operator]. The same scaling has also been found [parker2019universal] for the exactly solvable Sachdev–Ye–Kitaev model [Sachdev_1993_GaplessSpinFluid, Sachdev_1993_GaplessSpinFluid, Maldacena_2016_RemarksSYK] that describes fermions with random infinite-range couplings.
Here we provide the confirmation of UOGH for locally interacting nonintegrable fermionic systems in two and three dimensions. Our results clearly confirm the linear scaling for the 2D and 3D $t$-$V$ model, see the insets in Fig. 2 (b), (c), as well as for the 2D Hubbard model, see the inset in Fig. 2 (e).
Remarkably, we also observe the approximately linear scaling for one-dimensional $t$-$V$ and Hubbard models that are integrable. We leave the discussion of the origin and implications of this fact for further work.
Applying the recursion method.
The recursion method strengthened by the UOGH has proved highly useful in accessing dynamics at times beyond $t^{\ast}$ [parker2019universal, Uskov_2024_Quantum, Teretenkov_2025_Pseudomode, shirokov2025quench]. The method as such is nothing else but solving coupled Heisenberg equations in the Lanczos basis. When specified for finding the autocorrelation function, the method amounts to solving coupled differential equations,
$\partial_{t}\varphi_{n}(t)=-b_{n+1}\varphi_{n+1}(t)+b_{n}\varphi_{n-1}(t),\qquad n=0,1,2,\dots$ (9)
with the convention $\varphi_{-1}(t)=0$ and initial condition $\varphi_{n}(0)=\delta_{n\,0}$. The autocorrelation function is then found from $C(t)=\varphi_{0}(t)$.
However, the number $n_{\rm max}$ of Lanczos coefficients that can be explicitly computed is necessarily limited, since the size of nested commutators $H,J]^{(n)}$ grows exponentially with $n$. If we were to truncate the infinite system of equations at $n=n_{\rm max}$, the resulting approximation would be too rough to access the long-time dynamics of $C(t)$. Fortunately, this limitation can be overcome by extrapolating the unknown Lanczos coefficients for $n>n_{\rm max}$ according to the UOGH (8). This approach has already proven successful for strongly-correlated spin systems [parker2019universal, Uskov_2024_Quantum, Teretenkov_2025_Pseudomode, shirokov2025quench]. Here we apply it to fermionic systems.
Specifically, we employ the linear fit (8) to extrapolate the Lanczos coefficients for $n>n_{\text{max}}$, while for $n\leq n_{\text{max}}$ the exact values of $b_{n}$ are used. The resulting hybrid set of coefficients is then substituted into eq. (9). In practice, the system (9) is truncated at a finite value $K\gg n_{\text{max}}$, chosen such that the dynamics in the chosen time window is unaffected by the truncation. In all calculations presented here, we use $K=1000$.
We note that in the 3D Hubbard model the linear asymptotics of the Lanczos coefficients is not obvious, in contract to other models, see Fig. 2. We suggest that this is simply because of the small number of Lanczos coefficients we are able to compute. We, however, assume that the UOGH holds also in this case, and proceed in the same way as in other models.
In Fig. 2 we show the time evolution of the current autocorrelation function (4) for the models (1) and (2) in one, two, and three spatial dimensions. Results are shown for two representative values of the interaction strength in the intermediate-to-strong coupling regime, $V=2$ and $V=4$. In one dimension, our results are benchmarked against exact diagonalization. In all dimensions, they are additionally benchmarked by rigorous upper and lower bounds (7). The results demonstrate that the recursion method is able to compute the autocorrelation function up to thermalization times, thereby capturing all relevant dynamical time scales.
Figure 3: Diffusion constant in the infinite-temperature limit at half-filling, $\langle n_{i\sigma}\rangle=1/2$, as a function of the interaction strength. Top row: spinless-fermion $t$-$V$ model (1). Bottom row: Hubbard model (2). Solid blue – the results of the recursion method, with the width of the curve indicating the estimated uncertainty. Dashed red – the $\kappa/V^{2}$ fit. For the 2D Hubbard model, this fit coincides with the perturbative result of ref[Kovacevic_2025_Toward], see [supplement] for details.
Transport.
The knowledge of current-current autocorrelation function unlocks transport properties of the system through the Kubo linear response theory [Mahan_2013_Many]. In particular, the direct-current (dc) conductivity, $\sigma_{\rm dc}$, and the charge diffusion constant, $D$, read $\sigma_{\rm dc}=\beta\,{\cal V}^{-1}\,{\rm Re}\int_{0}^{\infty}dt\langle J(t)J\rangle_{\beta}$ and $D=\delta N_{\rm tot}^{-2}\,{\rm Re}\int_{0}^{\infty}dt\langle J(t)J\rangle_{\beta}$ respectively. Here $\beta=1/T$ is the inverse temperature, $\cal V$} is the number of lattice sites, $\langle\dots\rangle_{\beta}$ is the thermal average, $N_{\rm tot}=\sum_{i,\sigma}n_{i\sigma}$ is the total number of fermions and $\delta N_{\rm tot}^{2}=\langle N_{\rm tot}^{2}\rangle_{\beta}-\langle N_{\rm tot}\rangle_{\beta}^{2}$. In fact, the integral over the current-current autocorrelation function can be effectively extracted directly from the Lanczos coefficients [Joslin_1986_Calculation, Wang_2024_Diffusion, Bartsch_2024_Estimation], which facilitates transport calculations [Wang_2024_Diffusion, Bartsch_2024_Estimation, Uskov_2024_Quantum]. In the high-temperature limit, the diffusion constant saturates at a finite value, while the dc conductivity scales linearly with the inverse temperature,
$\sigma_{\rm dc}\simeq\beta\,{\cal V}^{-1}\,\left(\left.\delta N_{\rm tot}^{2}\,D\right|_{\beta=0}\right).$ (10)
In Fig. 3, we present the infinite-temperature diffusion constants for two- and three-dimensional $t$-$V$ and Hubbard models as functions of the interaction strength $V$ (calculations are detailed in [supplement]). Perturbation theory predicts $D\sim 1/V^{2}$ at small $V$ [Kiely_2021_Transport, Vucicevic_2023_Charge]. Surprisingly, we find that this scaling extends to quite large values of $V$, far beyond the expected range of applicability of the perturbation theory.
There exists a compelling experimental [Brown_2019_Bad] and theoretical [Brown_2019_Bad, Mousatov_2019_bad, Vranic_2020_Charge, Kiely_2021_Transport, Vucicevic_2023_Charge] evidence that the linear scaling of conductivity with $\beta$, as expressed in Eq. (10), remains valid down to relatively low temperatures on the order of 1 or even less. Combined with this observation, our calculations yield results for the dc conductivity across a wide range of interaction strengths and temperatures.
Our result for the diffusion constant in the 2D Hubbard model is consistent with a recent tour de force perturbative calculation reported in Ref. [Kovacevic_2025_Toward]. Notably, the latter study has culminated several earlier attempts [Kiely_2021_Transport, Vucicevic_2023_Charge] that had missed certain relevant leading-order perturbative contributions. The intricacy of this perturbative calculation underscores the nontrivial nature of the hight-temperature quantum transport.
Majorana Strings
Mention similarity to Pauli Strings basis, but with the detail that now we need to canonicalize the strings and that it does not increases the memory but increases time. Finally, we discuss technical aspects of calculating the moments (5). To compute the $n$-fold nested commutator $H(V),J]^{(n)}$, we employ an operator basis of Majorana strings. A Majorana string in canonical form is $\Gamma=\gamma_{i_{1},\sigma_{1}}\gamma_{i_{2},\sigma_{2}}\dots\gamma_{i_{l},\sigma_{l}}$, where $\gamma_{2j-1,\sigma}=c_{j\sigma}+c_{j\sigma}^{\dagger}$ and $\gamma_{2j,\sigma}=-i(c_{j\sigma}-c_{j\sigma}^{\dagger})$. The canonical ordering is fixed by (i) strictly increasing Majorana indices $i_{1}<i_{2}<\dots<i_{l}$ and (ii) for each site, the spin-down operator is placed before the spin-up operator, i.e., $\gamma_{j,\downarrow}\gamma_{j,\uparrow}$. For a system of spinless (spinful) fermions on $\cal V$ sites, there are in total ${\cal D}=4^{\cal V}$ ( ${\cal D}=4^{2{\cal V}}$) canonical Majorana strings $\Gamma_{s}$, which form an operator basis $\mathcal{G}=\{\Gamma_{s}\}_{s=1}^{{\cal D}}$ orthogonal with respect to the Frobenius inner product introduced earlier. Any fermionic operator can be uniquely decomposed in this basis. Consequently, the $n$-fold commutator can be written as
$[H(V),J]^{(n)}=\sum_{k=1}^{\mathcal{N}_{n}}a_{k}\Gamma_{k},$ (11)
where $a_{k}$ are complex coefficients and $\mathcal{N}_{n}$ denotes the number of nonzero Majorana strings appearing at order $n$.
In computer memory, each Majorana string $\Gamma_{k}$ is stored simply as an ordered list of indices, while the coefficients $a_{k}$ may be represented using finite-precision arithmetic, infinite-precision arithmetic, or in symbolic form. Our numerical routines operate directly in the space of Majorana strings. In practice, we first express the Hamiltonians (1,2) and the observable (3) in the Majorana basis, and then iteratively construct the commutators $H,J]^{(n)}$ until memory constraints are reached.
Using infinite-precision arithmetic, the largest operators we obtained for the two-dimensional ttVV model (1) at $k=9$ contain $\mathcal{N}^{\mathrm{2D\text{-}tV}}_{9}\simeq 7\times 10^{8}$ Majorana strings and occupy approximately $43$ GB of memory, corresponding to an average of $\sim 60$ bytes per string. The number of terms $\mathcal{N}_{n}$ grows exponentially with $n$. For example, in the three-dimensional Hubbard model we observe $\mathcal{N}^{\mathrm{3D\text{-}Hub}}_{n}\sim 10^{n}$, making each subsequent moment roughly an order of magnitude more expensive to compute than the previous one.
Although our framework allows us to keep the coefficients $a_{k}$ in symbolic form and thereby construct symbolic surrogates of the operators $H(V),J]^{(n)}$, this is not required to obtain symbolic expressions for the moments $\mu_{2n}(V)$. Since $\mu_{2n}(V)$ is a polynomial in $V$ of degree $2n$, it is sufficient to compute the moments at $n+1$ distinct values of $V$ using infinite-precision arithmetic and then reconstruct the full polynomial via interpolation. This strategy trades computational time for memory and enables access to higher-order moments in memory-limited settings.
Summary and outlook
We have implemented the symbolic recursion method for strongly interacting fermions on two and three dimensional lattices, focusing on $t$-$V$ and Hubbard models. The method yields moments and Lanczos coefficients of dynamic autocorrelation functions as functions of model parameters. We have verified that the universal operator growth hypothesis (UOGH) holds for locally interacting fermionic systems. It has been shown that recursion method supplemented by UOGH describes the dynamics for all time scales up to thermalization times. This gives access to transport properties. We have computed the diffusion constant in the limit of infinite temperature and the dc conductivity at temperatures above the model bandwidths.
Our work opens the door to further developments of operator-based methods in fermionic systems. A direct extension of the present results is the computation of higher-order moments by exploiting symmetries in operator space [teng2025leveraging], in the present work we made use only of translational symmetry. Another promising direction is the application of the recursion method to quench dynamics [shirokov2025quench] and to finite-temperature calculations [angelinos2025temperature].
We have also demonstrated the effectiveness of a symbolic computational workflow. In the future, one may obtain families of symbolic expressions depending on multiple parameters, such as the chemical potential, next-nearest-neighbour hopping amplitudes, or interaction strengths, for a variety of lattice geometries. A single upfront computational investment can then yield compact symbolic representations that allow for efficient evaluation of dynamical and transport quantities, across wide regions of parameter space.
Acknowledgements.
Acknowledgments.
The authors are grateful to I. Shirokov for useful discussions.