### v2

# supplement.pdf

Abstract
We present a symbolic implementation of the recursion method for the real-time dynamics of strongly correlated fermions on one-, two- and three-dimensional lattices. Focusing on two paradigmatic models – interacting spinless fermions and the Hubbard model – we find that the behavior of Lanczos coefficients is compatible with the universal operator growth hypothesis, exhibiting an approximately linear growth. Leveraging symbolically computed Lanczos coefficients and their asymptotic behavior, we access infinite-temperature autocorrelation functions up to times sufficient for thermalization to occur, directly in the thermodynamic limit. This, in turn, unlocks transport properties. In particular, we compute with high precision the charge diffusion constant over a broad range of interaction strengths, $V$. Surprisingly, the results are well described by a simple functional dependence $\sim 1/V^{2}$, which holds universally both for small and large $V$. Our results highlight the promise of a symbolic computational paradigm in which the most computationally expensive step is performed once, yielding reusable symbolic output that enables efficient evaluation of physical observables for arbitrary model parameters.
Dynamics and transport in strongly correlated electron systems, such as the spinless-fermion $t$-$V$ model and the Hubbard model, have been at the forefront of condensed-matter physics for decades [46, 24, 38, 79, 80, 18, 58, 34, 44]. Accurate simulation of these models remains one of the central challenges of modern theoretical physics. Approaches based on the direct diagonalization of Hamiltonians are fundamentally limited by the exponential growth of the many-body Hilbert space with the system size [94, 39]. In one dimension, remarkable progress has been achieved using integrability-based techniques [65, 12] and tensor-network methods [92], to the point that a broad class of dynamical and transport problems, with certain reservations, can now be considered simulable in 1D. In higher dimensions the substantial progress has been achieved using a variety of analytical and numerical approaches [8, 45, 23, 63, 11, 73, 61, 91, 41, 82, 81, 89, 44, 54, 31]. In parallel, considerable efforts have focused on quantum simulation platforms, which aim to realize synthetic correlated many-body systems and probe their properties experimentally [71, 5, 85, 22, 2]. Despite impressive recent progress, such platforms remain at an early stage of development and would greatly benefit from reliable computational benchmarks capable of guiding both experiments and theory.
Figure 1: Reusable symbolic computation workflow: the computationally expensive symbolic evaluation of the $n$-fold commutators $H(V),A]^{(n)}$ is carried out once, while the resulting symbolic moments $\mu_{2n}(V)$ are subsequently reused multiple times to compute correlation functions $C(t)$ at specific values of the parameter $V$.
In this work, we develop a symbolic computational approach to the dynamics of strongly correlated fermionic systems based on the recursion method. While first introduced decades ago [55, 26, 35, 53, 86], the recursion method has seen renewed attention due to recent theoretical progress on operator growth [28, 16, 60, 7, 25, 78, 57, 33, 48, 62] and the increase of computational capabilities.
The recursion method has been successfully applied in a variety of settings, including interacting spin-$1/2$ systems [87, 96, 95, 83, 50, 49, 74] and higher-spin models [30]. Other methods to address quantum many-body dynamics in Heisenberg representation have also emerged [66, 68, 29, 72, 10, 4, 69], typically formulated within qubit operator spaces where the Pauli-string basis provides a natural framework.
Fermionic models pose additional challenges. In Monte Carlo methods, these challenges manifest themselves as the infamous sign problem. In operator-based approaches, difficulties arise from the anticommutation relations of fermionic operators on spatially separated sites.
We report the successful application of the advanced recursion method to interacting fermionic systems on one-, two-, and three-dimensional lattices. In one dimension, the recursion method is known to be less efficient than state-of-the-art approaches based on exact diagonalization or matrix product states [97, 83, 74]. For this reason, applying the recursion method in one dimension serves the sole purpose of cross-checking its results against those from exact diagonalization. In contrast, real-time dynamics in two and three spatial dimensions is widely believed to be challenging for established numerical methods. We demonstrate that the advanced recursion method is a powerful technique applicable in higher dimensions, capable of capturing the dynamics over the entire time domain – from short times to thermalization – in the strongly correlated regime and without finite-size effects.
The rest of the manuscript is organized as follows. In the next section, we introduce specific models under study. Then we discuss the autocorrelation function and its moments. After that, we briefly review the recursion method and the universal operator growth hypothesis (UOGH) [60], present Lanczos coefficients for interacting fermionic models and demonstrate that they are compatible with the UOGH. Subsequently, we present infinite-temperature autocorrelation functions obtained from the recursion method supplemented by the UOGH. Then we apply the recursion method to describe transport in the models under study. Specifically, we compute the diffusion constant and the conductivity for a broad range of interaction strengths. In the End Matter, we describe important details of the algorithm and software implementation. Further details are relegated to the Supplemental Material [1].
Models and observables.
We consider two fermionic models on hypercubic lattices. The first one is the spinless fermion $t$-$V$ model,
$H_{tV}$ $=-\,t_{\text{hop}}\sum_{\langle i,j\rangle}\left(c_{i}^{\dagger}c_{j}+\text{H.c.}\right)+V\sum_{\langle i,j\rangle}\Bigl(n_{i}-\tfrac{1}{2}\Bigr)\Bigl(n_{j}-\tfrac{1}{2}\Bigr).$ (1)
The second one is the Hubbard model,
$H_{\mathrm{Hub}}$ $=-t_{\text{hop}}\sum_{\langle i,j\rangle,\sigma}\left(c_{i\sigma}^{\dagger}c_{j\sigma}+\text{H.c.}\right)$
$~~~+V\sum_{i}\left(n_{i\uparrow}-\tfrac{1}{2}\right)\left(n_{i\downarrow}-\tfrac{1}{2}\right).$ (2)
Here $c^{\dagger}_{i\sigma}$ and $c_{i\sigma}$ are fermionic creation and annihilation operators on site $i$, with $\sigma=\uparrow,\downarrow$ being the spin index. In the spinless $t$-$V$ model the spin index is omitted. Fermionic operators satisfy the usual anticommutation relations, $\{c_{i\sigma},c_{j\tau}^{\dagger}\}=\delta_{ij}\,\delta_{\sigma\tau}$. The parameters $t_{\text{hop}}$ and $V$ refer to the hopping amplitude and interaction strength, respectively. We set $t_{\text{hop}}=1$ unless stated otherwise. The symbol $\langle i,j\rangle$ labels pairs of nearest–neighbour sites. We adopt the convention $\hbar=1$ and $k_{\rm B}=1$ throughout the paper. In one dimension both models are integrable [47, 76].
We choose the total current along a fixed lattice direction, $\hat{x}$, as an observable of interest. For the Hubbard model the current reads
$J=-\,i\sum_{i,\sigma}\left(c_{i\sigma}^{\dagger}c_{i+\hat{x},\sigma}-c_{i+\hat{x},\sigma}^{\dagger}c_{i\sigma}\right),$ (3)
where $i+\hat{x}$ denotes the nearest neighbour of site $i$ displaced by one lattice spacing along the chosen direction (in one dimension this reduces to $i+1$). For the $t$-$V$ model the current is given by the same formula with subscripts $\sigma$ dropped.
Figure 2: Dynamical infinite-temperature autocorrelation function (4). Top row: spinless-fermion $t$-$V$ model (1). Bottom row: Hubbard model (2). Solid lines: results of the recursion method in the thermodynamic limit for the two values of interaction strengths $V$. Gray dashed lines indicate upper and lower Taylor polynomial bounds (7). Green dots in panels (a) and (d) display the exact diagonalization results for finite periodic chains (of lengths 1212 and $7$, respectively) evaluated for times small enough to be representative for the thermodynamic limit. Insets show the Lanczos coefficients and their linear extrapolation.
Autocorrelation function and moments.
We introduce an inner product in the space of operators according to $(A|B)\equiv\mathrm{tr}(A^{\dagger}B)/\mathrm{dim}(\mathcal{H})$, where $A$ and $B$ are arbitrary operators and $\mathrm{dim}(\mathcal{H})$ is the (finite) Hilbert space dimension. The corresponding operator norm reads $\lVert A\rVert=\sqrt{(A|A)}$.
We wish to compute the dynamical infinite-temperature autocorrelation function of the current defined as
$C(t)\equiv\frac{(J(t)|J)}{\lVert J\rVert^{2}}.$ (4)
Here $J(t)=e^{it\mathcal{L}}J$ is the current operator (3) in the Heisenberg representation, with $\mathcal{L}\equiv[H,\cdot]$ being the Liouvillian superoperator acting on the space of operators.
Note that since the models we consider conserve the total number of fermions (in the Hubbard model – separately for each polarization), the equilibrium state of the system is characterized by average density of fermions, in addition to temperature. The above definition of the autocorrelation function corresponds to half-filling, i.e. to $\langle n_{j}\rangle=1/2$ and $\langle n_{j\sigma}\rangle=1/2$ for the $t$-$V$ and Hubbard models, respectively.
Moments of the autocorrelation function are defined as
$\mu_{2n}\equiv\frac{(\mathcal{L}^{2n}J|J)}{\lVert J\rVert^{2}}=\frac{(\mathcal{L}^{n}J|\mathcal{L}^{n}J)}{\lVert J\rVert^{2}}=(-1)^{n}\frac{d^{\mkern 1.0mu2n}}{dt^{2n}}C(t)\big|_{t=0}.$ (5)
Note that this definition implies $\mu_{0}=1$.
Importantly, moments are polynomials of Hamiltonian parameters. Specifically, for the Hamiltonians (1) and (2), the moments (5) are polynomials in $V$, $\mu_{2n}(V)=\sum_{k=1}^{n}c^{(n)}_{k}V^{2k}$. In addition, the fermionic commutation relations ensure that coefficients $c^{(n)}_{k}$ for the current operator (3) are rational numbers. In case of $t_{\text{hop}}\neq 1$, the moments scale as $\mu_{2n}(t_{\text{hop}},V)=(t_{\text{hop}})^{2n}\mu_{2n}(V/t_{\text{hop}})$.
The core routine of our code outputs the first $n_{\text{max}}$ moments as polynomials in the system parameters. The computation is performed directly in the thermodynamic limit. We have obtained the sets of coefficients $c^{(n)}_{k}$ [75] for the Hamiltonians (1) and (2) in one, two and three dimensions. For example, the first few moments of the one-dimensional ttVV model (1) read
$2\mu_{2}(V)=V^{2},$
$2\mu_{4}(V)=5V^{2}+V^{4},$
$2\mu_{6}(V)=70V^{2}+21V^{4}+V^{6}.$ (6)
Once the first $n_{\text{max}}$ moments are computed, one can extract rich information about $C(t)$. In particular, one can obtain upper and lower bounds on the autocorrelation function [64, 67, 17, 15]. The simplest bound reads [83]
$P_{4l\pm 2}(t)\leq C(t)\leq P_{4l}(t),$ (7)
where $P_{m}(t)$ are Taylor polynomials of order $m$ constructed from the moments of order $m$ or less [1]. The bounds (7) are extremely tight initially but rapidly diverge at later times. The scope of these bounds is ultimately limited by the convergence radius $t^{\ast}$ of the Taylor expansion, which is finite for 2D and 3D systems [60].
Although not the central result of the present work, the above bounds serve as a valuable benchmark for the early-time dynamics of $C(t)$, in some cases extending to relatively long times. We provide the moments required to construct these bounds in symbolic form [75], making them readily available for straightforward benchmarking of other numerical methods and quantum simulations of the $t$-$V$ and Hubbard models for arbitrary values of interaction strength.
Lanczos coefficients and UOGH.
[…]
In fact, Lanczos coefficients $b_{n}$ can be obtained directly from moments $\mu_{2n}$ [26]. We employ this relation reviewed in [1] to symbolically compute Lanczos coefficients as functions of the interaction strength for the $t$-$V$ and Hubbard models in one, two and three dimensions, with the results illustrated in Fig 2.
An important question is the asymptotic behavior of Lanczos coefficients at large $n$. It can be proven for locally interacting systems that $b_{n}$ can not grow with $n$ faster than linear (with an additional logarithmic correction for one-dimensional systems) [60]. The UOGH [60] asserts that for chaotic quantum many-body systems this bound is saturated, i.e
$b_{n}=\alpha n+\gamma+o(1)\qquad{\rm as}\qquad n\rightarrow\infty$ (8)
in two and three spatial dimensions. For integrable systems the scaling is not universal: while typically it is either $\sqrt{n}$ or $O(1)$ [60], linear scaling has also been observed [27, 13, 20, 6, 84, 36].
The UOGH has been previously probed in various spin-1/2 systems on one- [60, 59, 37, 25], two- [37, 83, 25] and three-dimensional [74] lattices, for higher-spin systems on one- and two-dimensional lattices [30], as well as for a three-site system of interacting bosons [14]. A rigorous proof of the UOGH has been obtained for a class of spin-1/2 systems on hypercubic lattices [21]. Finally, the UOGH has been confirmed [60] for the exactly solvable Sachdev–Ye–Kitaev model [70, 70, 52] that describes fermions with random infinite-range couplings.
Here we probe the UOGH for locally interacting nonintegrable fermionic systems in two and three dimensions. Our results are clearly consistent with the linear asymptotic scaling of $b_{n}$ for the 2D and 3D $t$-$V$ model, see the insets in Fig. 2 (b), (c), as well as for the 2D Hubbard model, see the inset in Fig. 2 (e).
For the 3D Hubbard model, the linear asymptotics of the Lanczos coefficients is not apparent from our data shown in Fig. 2(f), likely due to the limited number of coefficients we can compute. We therefore treat the asymptotic linear growth of $b_{n}$ in this model as a working assumption, which we adopt for the subsequent analysis.
Remarkably, we also observe the approximately linear scaling for the one-dimensional $t$-$V$ and Hubbard models, which is somewhat unexpected for Bethe-ansatz-integrable models. We leave the discussion of the origin and implications of this fact for further work.
Applying the recursion method.
The recursion method strengthened by the UOGH has proved highly useful in accessing dynamics at times beyond $t^{\ast}$ [60, 83, 78, 74]. The method as such is nothing else but solving coupled Heisenberg equations in the Lanczos basis. When specified for finding the autocorrelation function, the method amounts to solving coupled differential equations,
$\partial_{t}\varphi_{n}(t)=-b_{n+1}\varphi_{n+1}(t)+b_{n}\varphi_{n-1}(t),\qquad n=0,1,2,\dots$ (9)
with the convention $\varphi_{-1}(t)=0$ and initial condition $\varphi_{n}(0)=\delta_{n\,0}$. The autocorrelation function is then found from $C(t)=\varphi_{0}(t)$.
However, the number $n_{\rm max}$ of Lanczos coefficients that can be explicitly computed is necessarily limited, since the size of nested commutators $H,J]^{(n)}$ grows exponentially with $n$. If we were to truncate the infinite system of equations at $n=n_{\rm max}$, the resulting approximation would be too rough to access the long-time dynamics of $C(t)$. Fortunately, this limitation can be overcome by extrapolating the unknown Lanczos coefficients for $n>n_{\rm max}$ according to the UOGH (8). This approach has already proven successful for strongly-correlated spin systems [60, 83, 78, 74]. Here we apply it to fermionic systems.
Specifically, we employ the linear fit (8) to extrapolate the Lanczos coefficients for $n>n_{\text{max}}$, while for $n\leq n_{\text{max}}$ the exact values of $b_{n}$ are used. The resulting hybrid set of coefficients is then substituted into eq. (9). In practice, the system (9) is truncated at a finite value $K\gg n_{\text{max}}$, chosen such that the dynamics in the chosen time window is unaffected by the truncation. For all calculations reported here we use $K=1000$.
In Fig. 2 we show the time evolution of the current autocorrelation function (4) for the models (1) and (2) in one, two, and three spatial dimensions. Results are shown for two representative values of the interaction strength in the intermediate-to-strong coupling regime, $V=2$ and $V=4$. In one dimension, our results are benchmarked against exact diagonalization. In all dimensions, they are additionally benchmarked by upper and lower bounds (7). We find that the recursion method computes the autocorrelation function up to thermalization times, capturing all relevant dynamical timescales.
Figure 3: Diffusion constant at half-filling in the infinite-temperature limit as a function of the interaction strength. Top row: spinless-fermion $t$-$V$ model (1). Bottom row: Hubbard model (2). Solid blue – the results of the recursion method, with the width of the curve indicating the estimated uncertainty. Dashed red – the $\kappa/V^{2}$ fit. For the 2D Hubbard model, this fit is consistent with the perturbative result of ref[43], see [1] for details.
Transport.
The knowledge of current-current autocorrelation function unlocks transport properties of the system through the Kubo linear response theory [51]. In particular, the direct-current (dc) conductivity, $\sigma_{\rm dc}$, and the charge diffusion constant, $D$, read $\sigma_{\rm dc}=\beta\,{\cal V}^{-1}\,{\rm Re}\int_{0}^{\infty}dt\langle J(t)J\rangle_{\beta}$ and $D=\delta N_{\rm tot}^{-2}\,{\rm Re}\int_{0}^{\infty}dt\langle J(t)J\rangle_{\beta}$ respectively. Here $\beta=1/T$ is the inverse temperature, $\cal V$} is the number of lattice sites, $\langle\dots\rangle_{\beta}$ is the thermal average, $N_{\rm tot}=\sum_{i,\sigma}n_{i\sigma}$ is the total number of fermions and $\delta N_{\rm tot}^{2}=\langle N_{\rm tot}^{2}\rangle_{\beta}-\langle N_{\rm tot}\rangle_{\beta}^{2}$. At infinite temperature, $\langle J(t)J\rangle_{\beta}=\lVert J\rVert^{2}C(t)$ with $C(t)$ defined by eq.(4). In fact, the integral over the current-current autocorrelation function can be effectively extracted directly from the Lanczos coefficients [40, 93, 9], which facilitates transport calculations [93, 9, 83]. In the high-temperature limit, the diffusion constant saturates at a finite value, while the dc conductivity scales linearly with the inverse temperature,
$\sigma_{\rm dc}\simeq\beta\,{\cal V}^{-1}\,\left(\left.\delta N_{\rm tot}^{2}\,D\right|_{\beta=0}\right).$ (10)
In Fig. 3, we present the infinite-temperature diffusion constants for two- and three-dimensional $t$-$V$ and Hubbard models as functions of the interaction strength $V$ (calculations are detailed in [1]). Perturbation theory predicts $D\sim 1/V^{2}$ at small $V$ [42, 90]. Surprisingly, we find that this scaling extends to quite large values of $V$, far beyond the expected range of applicability of the perturbation theory.
There exists a compelling experimental [19] and theoretical [19, 56, 88, 42, 90] evidence that the linear scaling of conductivity with $\beta$, as expressed in Eq. (10), remains valid down to relatively low temperatures on the order of 1 or even less. Combined with this observation, our calculations yield results for the dc conductivity across a wide range of interaction strengths and temperatures.
Our result for the diffusion constant in the 2D Hubbard model is consistent with a recent tour de force perturbative calculation reported in Ref. [43]. Notably, the latter study has culminated several earlier attempts [42, 90] that had missed certain relevant leading-order perturbative contributions. The intricacy of this perturbative calculation underscores the nontrivial nature of the hight-temperature quantum transport.
Summary and outlook
We have implemented the symbolic recursion method for strongly interacting fermions on two and three dimensional lattices, focusing on $t$-$V$ and Hubbard models. The method yields moments and Lanczos coefficients of dynamic autocorrelation functions as symbolic computer formulas involving model parameters. We have verified that the Lanczos coefficients for the above locally interacting fermionic models are generally consistent with the universal operator growth hypothesis (UOGH). It has been demonstrated that the recursion method supplemented by UOGH is capable of computing dynamical correlation functions up to thermalization times and accurately predicting transport properties. As an illustration, we have computed the diffusion constant in the limit of infinite temperature and the dc conductivity at temperatures above the model bandwidths. Quite unexpectedly, we have found that the diffusion constant obeys a simple scaling law over a broad range of interaction strengths.
Our work paves the way for further developments of operator-based methods in fermionic systems. Immediate extensions include reducing computational complexity by exploiting additional symmetries [77] – beyond the translational symmetry employed here – as well as applying the method to quench dynamics [48, 74] and extending it to finite temperatures [3].
Our work also demonstrates the remarkable effectiveness of a symbolic computational workflow: a single upfront computational investment yields a reusable symbolic formula that allows for subsequent efficient evaluation of dynamical and transport quantities for arbitrary model parameters. This paradigm, bridging analytic insights with computational scalability, opens new pathways for systematically exploring correlated quantum matter beyond the reach of conventional numerical methods.
Acknowledgements.
Acknowledgments.
The authors are grateful to I. Shirokov for useful discussions.
Data availability.
The core data required to reproduce our results – the symbolic moments (5) – is publicly available [75].
I End Matter
Here, we discuss technical aspects of computing commutators in the basis of Majorana strings and the moments (5). To compute the $n$-fold nested commutator $H(V),J]^{(n)}$, we employ an operator basis of Majorana strings. A Majorana string is a product of Majorana operators in the canonical order, $\Gamma=\gamma_{i_{1},\sigma_{1}}\gamma_{i_{2},\sigma_{2}}\dots\gamma_{i_{l},\sigma_{l}}$, where $\gamma_{2j-1,\sigma}=c_{j\sigma}+c_{j\sigma}^{\dagger}$ and $\gamma_{2j,\sigma}=-i(c_{j\sigma}-c_{j\sigma}^{\dagger})$. The canonical order implies that (i) the site labels are arranged in the strictly increasing order, $i_{1}<i_{2}<\dots<i_{l}$ and (ii) for each site, the spin-down operator is placed before the spin-up operator, i.e., $\gamma_{j,\downarrow}\gamma_{j,\uparrow}$. For a system of spinless (spinful) fermions on $\cal V$ sites, there are in total ${\cal D}=4^{\cal V}$ ( ${\cal D}=4^{2{\cal V}}$) canonical Majorana strings $\Gamma_{s}$, which form an operator basis $\mathcal{G}=\{\Gamma_{s}\}_{s=1}^{{\cal D}}$ orthogonal with respect to the inner product introduced earlier.
Any fermionic operator can be uniquely decomposed in this basis of Majorana strings. Consequently, the $n$-fold commutator can be written as
$[H(V),J]^{(n)}=\sum_{k=1}^{\mathcal{N}_{n}}a_{k}\Gamma_{k},$ (11)
where $a_{k}$ are complex coefficients and $\mathcal{N}_{n}$ denotes the number of nonzero Majorana strings appearing at order $n$.
In computer memory, each Majorana string $\Gamma_{k}$ is stored as an ordered list of indices, while the coefficients $a_{k}$ may be represented using finite-precision arithmetic, infinite-precision arithmetic, or in symbolic form. Our numerical routines operate directly in the space of Majorana strings. In practice, we first express the Hamiltonians (1,2) and the observable (3) in the Majorana basis, and then iteratively construct the commutators $H,J]^{(n)}$ until memory constraints are reached. Note that managing Majorana strings for fermions is strongly analogous to managing Pauli strings for qubits but requires the additional step of maintaining the canonical operator ordering, which incurs extra computational overhead. This extra step increases computational time as each string must be ordered, but does not increase the memory required.
Using infinite-precision arithmetic, the largest operators we obtained for the two-dimensional ttVV model (1) at $k=9$ contain $\mathcal{N}^{\mathrm{2D\text{-}tV}}_{9}\simeq 7\times 10^{8}$ Majorana strings and occupy approximately $43$ GB of memory, corresponding to an average of $\sim 60$ bytes per string. The number of terms $\mathcal{N}_{n}$ grows exponentially with $n$, see Fig. 4. For example, in the three-dimensional Hubbard model we observe $\mathcal{N}^{\mathrm{3D\text{-}Hub}}_{n}\sim 10^{n}$, making each subsequent moment roughly an order of magnitude more expensive to compute than the previous one.
Figure 4: Number of terms $\mathcal{N}_{n}$ in the $n$-th nested commutator (11) as a function of the commutator order $n$ for the ttVV and Hubbard models in one, two, and three dimensions.
Although our framework allows us to keep the coefficients $a_{k}$ in symbolic form and thereby construct symbolic surrogates [32, 68] of the operators $H(V),J]^{(n)}$, this is not required to obtain symbolic expressions for the moments $\mu_{2n}(V)$. Since $\mu_{2n}(V)$ is a polynomial in $V$ of degree $2n$, it is sufficient to compute the moments at $2n+1$ distinct values of $V$ using infinite-precision arithmetic and then reconstruct the full polynomial via interpolation. This strategy trades computational time for memory and enables access to higher-order moments in memory-limited settings.
Symbolic Majorana strings are also utilized in a recently proposed Majorana propagation method [54]. This method mostly focuses on simulating fermionic quantum circuits, but can also be applied to simulations of real-time dynamics by approximating the actual Heisenberg evolution by discrete Trotter steps. Recent applications of the method to small finite two-dimensional Hubbard models show that Majorana propagation is efficient in the case of small interaction strength $V$ [31], but can face limitations for larger values of $V$ [31, 2].